OpenStudy (anonymous):

2x + 3y = 6 5x + 2y = 4 Which of the following equations could be the result of multiplication and addition to eliminate a variable in the system of equations?

OpenStudy (anonymous):

multiply eq.1 by 2 and eq. 2 by 3 and then subtract

OpenStudy (anonymous):

y wil get eliminated

OpenStudy (anonymous):

11y=22

OpenStudy (johnweldon1993):

actually i believe you would multiply the equation by -2 and then the second one by 3 this will give you -4x + -6y = -12 15x + 6y = 12 because when you combine them....you add...THIS was -6y and +6y cancel out so you would have -4x + 15x is? 11x -6y + 6y is? 0 -12 + 12 is? 0 so you have 11x = 0 x would equal 0.... you can do the same thing to solve for y multiply the top one by -5 and the bottom one by 2...this will make the x's cancel -10x - 15y = -30 10x + 4y = 8 now when you combine them you have -10x + 10x is? 0 -15y + 4y is? -11y -30 + 8 is? -22 so you have -11y = -22 divide both sides by -11 y = 2 now check 2(0) + 3(2) = 6 0 + 6 = 6 yes 5(0) + 2(2) = 4 0 + 4 = 4 yes so those are your answers :)

OpenStudy (anonymous):

so y =2 then find x

OpenStudy (anonymous):

5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is ___.

OpenStudy (johnweldon1993):

so you would do the same thing...just solve for y how do we get y by itself? eliminate x multiply the top equation by -1 -5x - y = -6 5x + 3y = -4 now when you add the two...the x's cancel out so we have 2y = -10 divide both sides by 2 y = -5 this would be the y coordinate :)

OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown?